At age 9, the average weight (21.3 kg) and the average height (124.5 cm) for both boys and girls are exactly the same. A randa. estimate the mean difference in height between boys and girls with a 95% confidence

b. does this interval support the given claim
Answer #1

solution Hele ni= 50, 71 = 123.8, 7 =98 n2 = 50, T2 =122-8, 3= 120 At as%. , Za/2=1.96 The 95% (-I for lilla es = (1238 - 122